Part 1 of 1 - Just Baked Just Baked cupcakes sold at retail stores for prices ranging from about $2.30 each for simpler cupcaWhat is Just Bakeds underage? O A. $2.00 OB. $2.75 OC. $2.25 OD. $.50 Reset SelectionWhat is Just Bakeds overage? O A. $2.75 OB. $2.00 O C. $2.25 OD. $.50 Reset SelectionWhat is Just Bakeds critical ratio? ΟΑ. .91 Ο Β. .81 OC. .61 Ο D. .71 Reset SelectionHow many cupcakes should they bake for each day? O A. 982 OB. 1,382 O C. 1,582 OD. 1,782 Reset Selection
Answer #1

The correct answer is option C.$ 2.25

Underage=Cost of understocking=Cu=Selling price-Cost price=2.75-0.5=$ 2.25

The correct answer is option D.$ 0.50

Overage=Cost of overstocking=Co=Cost price-Salvage value=0.5-0=$ 0.50

The correct answer is option B.0.81

Critical ratio=Cu/(Cu+Co)=2.25/(2.25+0.5)=0.8181=0.81

The correct answer is option B.1382

We have,

Optimal CSL(cycle service level)=critical ratio=0.8181

We have in normal distribution,

z=x-mean/std dev

Using excel function,we calculate z as

z=NORMSINV(CSL)=NORMSINV(0.8181)

  z=0.9081

solving for x ,we get,

0.9081=x-1200/200

x=1381.6=1382 cupcakes